OpenStudy (anonymous):

A depth finder shows that the water in a certain place is 620 feet deep. The diffidence between d, the actual depth of the water, and the reading is |d-620| and must be less than or equal to 0.05d. Find the minimum and maximum values of d, to the nearest tenth of a foot.

OpenStudy (anonymous):

is it "less than or equal to .05d" or "less than or equal to .05"?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[|d-320|\leq .05d\] means either \[d-320\leq .05d\] and \[d-320\geq -.05d\] so we have to solve these two seperately

OpenStudy (anonymous):

\[d-320\leq .05d\] \[.95d\leq 320\] \[d\leq 320\div .95\] \[d\leq 336.84\]

OpenStudy (anonymous):

that is the first one can you do the second? similar os is very slow today

OpenStudy (anonymous):

@satellite73 - your method is correct, but I think that you made a mistake by using |d-320| instead of |d-620|